# Saddle Hierarchy in Dense Associative Memory

### v1

@@ 2.2 Teacher-student setting
Among the extensive research on the properties of artificial neural networks (NNs) from the perspective of statistical mechanics [2, 3, 4, 14, 15, 16, 17, 18, 29, 62, 66, 67, 68, 69], there have been many studies of simple RBMs trained by maximum likelihood estimation (see Eq. 4) [70, 71, 72, 73, 74, 75] or by averaging samples from the posterior distribution of the weights $\mathbf{J}$ given some observed data $\mathcal{D}$. The latter approach has notably been used to characterize the fundamental limits of RBM learning [76, 77, 78, 79, 80, 81, 82, 83, 84] in the teacher-student setting where the data used to train $\mathbf{J}$ is sampled from another RBM with planted weights $\mathbf{J}^{*}$ [85, 78, 79, 86]. This teacher-student setting can also be used to study our DAM. In this scenario, a teacher DAM with weights $\mathbf{w}^{*}$ and $\mathbf{p}^{*}$ generates a large amount $M=\alpha N$ of noisy data $\mathcal{D}=\left\{\mathbf{x}^{c},y^{c}\right\}_{c=1}^{M}$ and feeds them to a student DAM, which then trains its weights $\mathbf{w}$ and $\mathbf{p}$ by averaging samples from the posterior distribution
@@ 3.1 Saddle-point equations
In Appendix B, we show that the stationary points of the negative log-likelihood loss (Eq. 4) satisfy
@@
for all $1\leq\mu\leq P$ and $0\leq\gamma\leq P$, where $m^{\mu_{*}0}\left(\mathbf{x}^{*},\mathbf{w}\right)=\frac{1}{\beta}\log\left[\Omega_{N}\left(\beta\right)/\Omega_{N}\left(0\right)\right]$ and the normalization constant $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ is defined in Appendix F.
@@
whose extremizer $\mathbf{m}=\left\{m^{\mu_{*}\mu}\right\}_{1\leq\mu\leq P}^{1\leq\mu_{*}\leq P^{*}}$ can be interpreted as the $N\rightarrow\infty$ limit of the expected value of the teacher-student overlaps $\mathbf{m}(\mathbf{w}^{*},\mathbf{w})$. We show the derivation of the variational principle in Appendix D together with an explicit expression for the trial function $f(\mathbf{m},\hat{\mathbf{m}},\mathbf{p})$. If we assume that there are $P^{*}\ll N$ teacher memories and that the priors on $\mathbf{w}^{*\mu_{*}}$ and $\mathbf{g}^{*}$ are uniform like those of the student, we find that the expected teacher-student overlaps must satisfy the saddle-point equations
@@
If we instead clamp the teacher weights $\mathbf{w}^{*\mu_{*}}$ and $\mathbf{g}^{*\mu_{*}}$ to fixed patterns $\mathbf{x}^{*\mu_{*}}$ and their corresponding soft labels $\mathbf{q}^{*\mu_{*}}$ [64], respectively, to mimic a more general distribution for the data, then Eqs. (3.1) become (see Appendix E)
@@
Let us explain this limit step by step. For any $\upsilon$ and $\varrho$, the $M$ examples $\{\mathbf{x}^{c}\}_{c=1}^{M}$ provided to the student are corrupted versions of the teacher patterns $\mathbf{x}^{*}$ with a noise level of $1/\upsilon$. Therefore, in the limit of $\upsilon\rightarrow\infty$ with finite $\varrho$, each example $\mathbf{x}^{c}$ provided to the student is one of the original patterns $\mathbf{x}^{*}$, as in Eqs. (3.1). However, even in this case, the empirical distribution of the examples deviates from that of $\mathbf{x}^{*}$, of which it is merely a bootstrap sample. At fixed $P^{*}$ and $P$, this mismatch disturbs the accurate learning of $\mathbf{x}^{*}$ when $\varrho=\frac{M}{P^{*}N}$ is finite, or equivalently the expected number of repetitions $M/P^{*}$ of each pattern is not sufficiently large compared to $N$ (see Eq. 3.1). This influence progressively weakens as $\varrho$ grows and the $\varsigma$ function approaches $1$, reflecting the convergence of the empirical distribution of the examples to that of the teacher patterns.
In the limit of $\varrho\rightarrow\infty$ with finite $\upsilon$, Eq. (3.1) is still very similar to Eq. (3.1). The only difference between them is that Eq. (3.1) has $\beta_{\text{eff}}$ instead of $\beta$ in the argument of $\sigma$ and in the denominator of the definition of $m^{\mu_{*}0}$. Using $\varsigma$ as a shorthand for $\varsigma\left(2\upsilon\right)$, we find that the fixed points of Eqs. (3.1) with $\varrho\rightarrow\infty$ are related, through the identification made by Eq. (11), to the stationary points of the effective loss
@@
What distinguishes this equation from the standard loss (Eq. 4) is that $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x},y\big{|}\mathbf{w},\mathbf{p}\right)$ has $\beta_{\text{eff}}=\varsigma\beta$ in the argument of the exponential function instead of $\beta$. As a consequence, $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x},y\big{|}\mathbf{w},\mathbf{p}\right)$ is not a probability distribution unless $\varsigma=1$, in the limit of $\upsilon\rightarrow\infty$ (see Appendix C). A value of $\varsigma$ less than one in the effective loss (Eq. 12) is reminiscent of the presence of noise in the data generation process, so we propose to use it as a regularizer for the weights. We discuss this point in more detail in Section 4.1.
@@ 3.2 Saddle-point hierarchy
As shown in [8], the loss landscape of any NN with unconstrained weights contains the stationary points of narrower NNs with the same architecture. In Appendix G, we show that this result also applies to the teacher-student setting with $\varrho\rightarrow\infty$ and any non-zero $\upsilon$. To be more precise, we show that, if the parameters $\bar{x}^{\text{fixed},\mu}_{i}$, $\bar{p}^{\text{fixed},\gamma}_{y}$, $m^{\text{fixed},\mu_{*}\gamma}$, $p^{\text{fixed},\gamma}_{y}$ with hidden unit prior $p_{\mathbf{h}}^{\text{given}}\left(\gamma\right)$ are a fixed point of Eqs. (3.1) with $P$ hidden units, then the duplicated parameters
@@
are a fixed point of the same saddle-point equations with $P+R\in\left\{P,...,2P\right\}$ hidden units (See Appendix G for a detailed derivation). In other words, we duplicate some of the weights solving Eqs. (3.1) to construct a fixed point for a wider network. In that sense, wide DAMs contain the fixed points of narrower DAMs. In particular (see Section 3.1), this property also holds for the stationary points of both the standard loss (Eq. 4) and the effective loss (Eq. 12).
The saddle-point equations (Eq. 3.1) with duplicated order parameters (Eqs. 3.2) are invariant to the permutation of any hidden unit $\gamma\in\left\{1,...,R\right\}$ and its duplicate $\gamma+P$. This kind of symmetry can be spontaneously broken if it leads to a higher free entropy (see Eqs. 7 and 22). However, symmetry-breaking transitions can be prohibitively slow when the symmetric state is stable to local perturbations [89]. Interestingly, the DAM introduced in [10] quickly undergoes many successive permutation symmetry-breaking bifurcations during training [11]. This observation and additional empirical evidence suggest that the symmetric states are unstable, or in other words that they are saddles, which was verified analytically for DAMs with only two data points to memorize [11]. In Appendix G, we prove that, if $\beta$ is large enough, Eqs. (3.2) is a saddle, which is a major step toward explaining why permutation symmetry-breaking transitions are relatively fast in DAMs trained with a large number of data points. We call this result the saddle-point hierarchy principle.
@@ 4 Empirical results
In this Section, we use our theoretical results to improve training, and we show empirically that our DAM learns interpretable solutions to both supervised and unsupervised classification problems. We perform all our numerical experiments on the MNIST dataset of handwritten digits [65]. The code and hyperparameter values used in our numerical experiments are available at this public repository [38].
@@ 4.1 Learning by minimizing the effective loss
As the number of hidden units $P$ increases, standard maximum likelihood estimation with fixed inverse temperature $\beta$ becomes progressively less apt to train our DAM to its full potential. At high $P$ and $\beta$, many memories stay stuck in noisy states that do not contribute to classification, which is wasteful. Using a lower $\beta$ helps the memories converge, but also reduces their resolution and diversity, possibly because it also lowers the DAM’s capacity [29]. This change is far from being only cosmetic. In fact, it comes with a gradual reduction in classification accuracy (see Table 1).
@@
Optimization algorithms with a parameter analogous to $\beta$ often converge to better solutions when this parameter is increased from a small value during optimization. For example, annealing schedules [90, 91] can be incorporated into minimization algorithms to help them find deeper local minima than they could otherwise. This observation suggests that our DAM would learn better memories if we increased $\beta$ during training. It is tempting to do so by SGD of Eq. (4) with respect to $\beta$, but it makes $\beta$ increase so quickly that many memories still stay stuck in noisy states.
@@ 4.2 Dense associative memory is interpretable, even in unsupervised classification
At test time, we observe that our DAM classifies approximately 98% of the test data points into the class $y=\operatorname*{argmax}_{y^{\prime}}\left\{p^{\mu}_{y^{\prime}}\right\}$ of the memory $\mathbf{w}^{\mu}$ to which they are the most similar. In other words, a 1-nearest neighbor classifier [92] conditioned on the memories $\mathbf{w}^{\mu}$ and their soft soft labels $\mathbf{p}^{\mu}/p_{\mathbf{h}}\left(\mu\right)$ approximates the classification of our model with 98% fidelity. This behavior is reminiscent of K & H’s DAM [10], where only a few memories participate in the classification of each data point. The 1-nearest neighbor classifier approximates DAM classification more faithfully for correctly classified data points (approximately 99% fidelity) than incorrectly classified data points (approximately 70% fidelity).
@@
As before, we do so using constrained SGD (see Appendix F) with the initialization and learning rate described in Appendix I. Equivalently, we can also view this algorithm as minimizing the combined loss
@@
where $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x}^{*\mu}|\mathbf{w},\mathbf{p}\right)=\sum_{y=0}^{C}\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x}^{*\mu},y|\mathbf{w},\mathbf{p}\right)$ (see Eq. 14). $\mathcal{L}_{\text{cond}}\left(\mathbf{w},\mathbf{p}\right)$ is called the minimum entropy regularization term in unsupervised machine learning [93]. Intuitively, it encourages the DAM to learn different class weights $p^{\gamma}_{y}$ for each class $y$, which is not possible by minimizing only $\mathcal{L}_{\text{margin}}\left(\mathbf{w},\mathbf{p}\right)$ because $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x}^{*\mu}|\mathbf{w},\mathbf{p}\right)$ does not depend on $y$. Exploiting this characteristic, we train our DAM on patches of MNIST digits [65] and find that it learns reasonable latent classes $y=\operatorname*{argmax}_{y^{\prime}}\left\{p^{\mu}_{y^{\prime}}\right\}$ for the memories $\mathbf{w}^{\mu}$ (see Fig. 4 and Figs. 7, 8, 9 and 10 of Appendix J). This approach could potentially be useful for feature extraction [94].
@@ 4.3 Fast training with splitting steepest descent
Machine learning models that experience the permutation symmetry-breaking bifurcations described in Section 3.2 during training—such as the DAM studied in [11], RBMs with binary units [81, 84] and Gaussian mixtures [62, 50, 51, 52, 91, 95, 96]—have characteristic tree-shaped learning dynamics. Intuitively, permutation symmetry-breaking transitions are points where model parameters differentiate from each other, so they correspond to the bifurcations of the tree. In the left panel of Fig. (5), we show that the learning dynamics of our DAM also follows a tree of permutation symmetry-breaking transitions.
@@
The saddle-point hierarchy principle introduced in Section 3.2 suggests the following idea to accelerate learning: train a relatively narrow DAM for cheap, then repeatedly duplicate (or “split”) some of the hidden units $\mu$, escape the corresponding saddle point and continue training. Intuitively, we expect to save a lot of computing resources by using relatively few hidden units at the start of training, when the learning dynamics follows a few branches close to the root of the learning dynamics tree. The splitting steepest descent algorithm introduced in [36] formalizes this idea in an efficient and theoretically motivated way. We implement a variant of this algorithm that takes the constraint $\mathbf{w}^{\mu}\in S^{N-1}$ into account, but is otherwise very similar to the original version (see Alg. 1), with fast splitting implemented as in [37]. The constraint $\mathbf{w}^{\mu}\in S^{N-1}$ only matters in steps 5 and 11, which are also arguably the most conceptually difficult parts of Alg. (1), so we explain them in detail in Appendix H. Fig. (5) shows that the learning dynamics tree of splitting steepest descent has a more sparsely populated trunk than that of SGD without splitting (see Fig. 5). In other words, by using a relatively small number $P$ of memories at the beginning of training, we reduce the total number of values $P\times T$ that they take over a period of time $T$, which is consistent with our intuition that it can save computational resources. The situation could have been different if using fewer memories slowed down training.
@@
Figure 5: Overlaps $m^{\mu_{*}\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)=\sum_{i=1}^{N}x^{*\mu_{*}}_{i}w^{\mu}_{i}$ between the first 1000 digits $\mathbf{x}^{*\mu_{*}}=\left\{x^{*\mu_{*}}_{i}\right\}_{i=1}^{N}$ of the MNIST training set [65] and the memories $\mathbf{w}^{\mu}=\left\{w^{\mu}_{i}\right\}_{i=1}^{N}$ of our dense associative memory (DAM) model while it is learning them. Each point is one of the high-dimensional magnetization vectors $m_{\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)=\left\{m^{\mu_{*}\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)\right\}_{\mu_{*}=1}^{P^{*}}$ projected onto a two-dimensional plane using the UMAP algorithm [97]. On the left, the $m_{\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)$ found during training without splitting steepest descent. On the right, the $m_{\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)$ found with splitting steepest descent (Alg. 1). In both cases, UMAP is trained on the $m_{\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)$ found without splitting steepest descent. The classes $y=\operatorname*{argmax}_{y^{\prime}}\left\{p^{\mu}_{y^{\prime}}\right\}$ of the memories $\mathbf{w}^{\mu}$ are color-coded in the legend. 7 and 1 are the most numerous classes, so they have the top two largest entries in $\sum_{\gamma=0}^{P}p^{\mu}_{y}=p_{\mathbf{q}}\left(y\right)$ (see Appendix I), which is why the memories are classified as either 7 or 1 at the beginning of training.
@@
14:return $\triangleright$ See Appendix H and [36, 37] for details about steps 5 and 11
We now establish a methodology that we will use to compare the training times of DAMs trained using the effective loss (Eq. 12) on MNIST [65] with and without splitting steepest descent. It is more interesting and meaningful to compare the training times of NNs with similar performance. Therefore, we pick hyperparameters such that DAMs trained with and without splitting have similar classification accuracy. For that purpose, we find the general region of hyperparameter space where DAMs trained without splitting steepest descent have the best classification accuracy. The accuracy does not change much in this region, so we pick generic hyperparameters inside. Next, we manually tune the hyperparameters of splitting steepest descent so that the resulting DAMs have comparable accuracy. Note that we did not find hyperparameters that make the accuracy with splitting steepest descent significantly better than without it.

### v2

@@ 2.2 Teacher-student setting
Among the extensive research on the properties of artificial Neural Networks (NNs) from the perspective of statistical mechanics (see [52] for a review), there have been many studies of simple RBMs trained by maximum likelihood estimation (a line of work pioneered in [53, 54, 55]) or by averaging samples from the posterior distribution of the weights $\mathbf{J}$ given some observed data $\mathcal{D}$. The latter approach has notably been used to characterize the fundamental limits of learning for many types of RBMs [56, 57, 58], with different priors [59, 60, 61] and architectures [62, 63], in the teacher-student setting [64, 65, 66] where the data used to train $\mathbf{J}$ is sampled from another RBM with planted weights $\mathbf{J}^{*}$. This teacher-student setting can also be used to study our DAM. In this scenario, a teacher DAM with weights $\mathbf{w}^{*}$ and $\mathbf{p}^{*}$ generates a large amount $M=\alpha N$ of noisy data $\mathcal{D}=\left\{\mathbf{x}^{c},y^{c}\right\}_{c=1}^{M}$ and feeds them to a student DAM, which then trains its weights $\mathbf{w}$ and $\mathbf{p}$ by averaging samples from the posterior distribution
@@
Table (2) of Appendix E summarizes, for quick reference, the most important symbols introduced in this Section and their meaning.
@@ 3.1 Saddle-point equations
In Appendix F, we show that the stationary points of the negative log-likelihood loss (Eq. 4) satisfy
@@
for all $1\leq\mu\leq P$ and $0\leq\gamma\leq P$, where $m^{\mu_{*}0}\left(\mathbf{x}^{*},\mathbf{w}\right)=\frac{1}{\beta}\log\left[\Omega_{N}\left(\beta\right)/\Omega_{N}\left(0\right)\right]$ and the normalization constant $\zeta^{\gamma}_{y}\left(\mathbf{\bar{p}};p_{\mathbf{h}}\right)$ is defined in Appendix C.
@@
whose extremizer $\mathbf{m}=\left\{m^{\mu_{*}\mu}\right\}_{1\leq\mu\leq P}^{1\leq\mu_{*}\leq P^{*}}$ can be interpreted as the $N\rightarrow\infty$ limit of the expected value of the teacher-student overlaps $\mathbf{m}(\mathbf{w}^{*},\mathbf{w})$. We show the derivation of the variational principle in Appendix G together with an explicit expression for the trial function $f(\mathbf{m},\hat{\mathbf{m}},\mathbf{p})$. If we assume that there are $P^{*}\ll N$ teacher memories and that the priors on $\mathbf{w}^{*\mu_{*}}$ and $\mathbf{g}^{*}$ are uniform like those of the student, we find that the expected teacher-student overlaps must satisfy the saddle-point equations
@@
If we instead clamp the teacher weights $\mathbf{w}^{*\mu_{*}}$ and $\mathbf{g}^{*\mu_{*}}$ to fixed patterns $\mathbf{x}^{*\mu_{*}}$ and their corresponding soft labels $\mathbf{q}^{*\mu_{*}}$ [49], respectively, to mimic a more general distribution for the data, then Eqs. (3.1) become (see Appendix H)
@@
Let us explain this limit step by step. For any $\upsilon$ and $\varrho$, the $M$ examples $\{\mathbf{x}^{c}\}_{c=1}^{M}$ given to the student are corrupted versions of the teacher patterns $\mathbf{x}^{*}$ with a noise level of $1/\upsilon$. Therefore, in the limit of $\upsilon\rightarrow\infty$ with finite $\varrho$, each example $\mathbf{x}^{c}$ is one of the original patterns $\mathbf{x}^{*}$, as in Eqs. (3.1). However, even in this case, the empirical distribution of the examples deviates from that of $\mathbf{x}^{*}$, of which it is merely a bootstrap sample. At fixed $P^{*}$ and $P$, this mismatch disturbs the accurate learning of $\mathbf{x}^{*}$ when $\varrho=\frac{M}{P^{*}N}$ is finite, or equivalently the expected number of repetitions $M/P^{*}$ of each pattern is not sufficiently large compared to $N$ (see Eq. 3.1). This influence progressively weakens as $\varrho$ grows and the $\varsigma$ function approaches $1$, reflecting the convergence of the empirical distribution of the examples to that of the teacher patterns.
In the limit of $\varrho\rightarrow\infty$ with finite $\upsilon$, Eq. (3.1) is still very similar to Eq. (3.1). In that case, the only difference between them is that Eq. (3.1) has $\beta_{\text{eff}}$ instead of $\beta$ in the argument of $\sigma$ and in the denominator of the definition of $m^{\mu_{*}0}$. Using $\varsigma$ as a shorthand for $\varsigma\left(2\upsilon\right)$, we find that the fixed points of Eqs. (3.1) with $\varrho\rightarrow\infty$ are related, through the identification made in Eq. (11), to the stationary points of the effective loss
@@
What distinguishes this equation from the standard loss (Eq. 4) is that $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x},y\big|\mathbf{w},\mathbf{p}\right)$ has $\beta_{\text{eff}}=\varsigma\beta$ in the argument of the exponential function instead of $\beta$. As a consequence, $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x},y\big|\mathbf{w},\mathbf{p}\right)$ is not a probability distribution unless $\varsigma=1$, in the limit of $\upsilon\rightarrow\infty$ (see Appendix B). A value of $\varsigma$ less than one in the effective loss (Eq. 12) is reminiscent of the presence of noise in the data generation process, so we propose to use it as a regularizer for the weights. We discuss this point in more detail in Section 4.1.
@@ 3.2 Saddle-point hierarchy
As shown in [8], the loss landscape of any NN with unconstrained weights contains the stationary points of narrower NNs with the same architecture. In Appendix I, we show that this result also applies to the teacher-student setting with $\varrho\rightarrow\infty$ and any nonzero $\upsilon$. To be more precise, we show that, if the parameters $\bar{x}^{\text{fixed},\mu}_{i}$, $\bar{p}^{\text{fixed},\gamma}_{y}$, $m^{\text{fixed},\mu_{*}\gamma}$, $p^{\text{fixed},\gamma}_{y}$ with hidden unit prior $p_{\mathbf{h}}^{\text{given}}\left(\gamma\right)$ are a fixed point of Eqs. (3.1) with $P$ hidden units, then the duplicated parameters
@@
are a fixed point of the same saddle-point equations with $P+R\in\left\{P,...,2P\right\}$ hidden units (See Appendix I for a detailed derivation and Fig. 2 for a illustrative example). In other words, we duplicate some of the weights solving Eqs. (3.1) to construct a fixed point for a wider network. In that sense, wide DAMs contain the fixed points of narrower DAMs. In particular (see Section 3.1), this property also holds for the stationary points of both the standard loss (Eq. 4) and the effective loss (Eq. 12).
$\bar{x}^{\text{fixed},\mu}_{i}$                  $\bar{x}^{\text{dupli},\mu}_{i}$
Figure 2: Illustration of the relationship between $\bar{x}^{\text{fixed},\mu}_{i}$ and $\bar{x}^{\text{dupli},\mu}_{i}$ stated in Eq. (3.2). The left panel represents the fixed point parameters $\bar{x}^{\text{fixed},\mu}_{i}$ of Eq. (3.1) with $P=10$, and the right panel represents the fixed point parameters $\bar{x}^{\text{dupli},\mu}_{i}$ of Eq. (3.1) with $P=15$. In this example, $R=5$, which means that the first 55 entries of $\bar{x}^{\text{fixed},\mu}_{i}$ are repeated twice in $\bar{x}^{\text{dupli},\mu}_{i}$, while the remaining ones are repeated only once. The first 1010 entries of $\bar{x}^{\text{dupli},\mu}_{i}$ are identical to $\bar{x}^{\text{fixed},\mu}_{i}$, and the dashed red lines highlights that the first 55 entries of $\bar{x}^{\text{fixed},\mu}_{i}$ are repeated a second time at the end of $\bar{x}^{\text{dupli},\mu}_{i}$.
The saddle-point equations (Eq. 3.1) with duplicated order parameters (Eqs. 3.2) are invariant to the permutation of any hidden unit $\gamma\in\left\{1,...,R\right\}$ and its duplicate $\gamma+P$. This kind of symmetry can be spontaneously broken if it leads to a higher free entropy (see Eqs. 7 and 26). However, symmetry-breaking transitions can be prohibitively slow when the symmetric state is stable to local perturbations [69]. Interestingly, the DAM introduced in [16] quickly undergoes many successive permutation symmetry-breaking bifurcations during training [24]. This observation and additional empirical evidence suggest that the symmetric states are unstable, or in other words that they are saddles, which was verified analytically for DAMs with only two data points to memorize [24]. In Appendix I, we prove that, if $\beta$ is large enough, Eqs. (3.2) is a saddle, which is a major step toward explaining why permutation symmetry breaking is relatively fast in DAMs trained with a large number of data points. We call this result the saddle-point hierarchy principle.
@@ 4 Empirical results
In this Section, we use our theoretical results to improve training, and we show empirically that our DAM learns interpretable solutions to both supervised and unsupervised classification problems. Unless explicitly stated otherwise, we perform our numerical experiments on the MNIST dataset of handwritten digits [50]. The code and hyperparameter values used in our numerical experiments are available in the following public repository [30].
@@ 4.1 Learning by minimizing the effective loss
As the number of hidden units $P$ increases, standard maximum likelihood estimation with fixed inverse temperature $\beta$ becomes progressively less apt to train our DAM to its full potential. At high $P$ and $\beta$, many memories stay stuck in noisy states that do not contribute to classification, which is wasteful (see Fig. 3). Using a lower $\beta$ helps the memories converge, but also reduces their resolution and diversity, possibly because it also lowers the DAM’s capacity [51]. This change is far from being only cosmetic. In fact, it comes with a gradual reduction in classification accuracy (see Table 1).
@@
Optimization algorithms with a parameter analogous to $\beta$ often converge to better solutions when this parameter is increased from a small value during optimization. For example, annealing schedules [70, 71] can be incorporated into minimization algorithms to help them find deeper local minima than they could otherwise. This observation suggests that our DAM would learn better memories if $\beta$ increased during training. It is tempting to do so by SGD of Eq. (4) with respect to $\beta$, but it makes $\beta$ increase so quickly that many memories still stay stuck in noisy states.
@@ 4.2 Dense associative memory is interpretable, even in unsupervised classification
At test time, we observe that our DAM classifies approximately 98% of the test data points into the class $y=\operatorname*{argmax}_{y^{\prime}}\left\{p^{\mu}_{y^{\prime}}\right\}$ of the memory $\mathbf{w}^{\mu}$ to which they are the most similar. In other words, a 1-nearest neighbor classifier [72] conditioned on the memories $\mathbf{w}^{\mu}$ and their soft labels $\mathbf{p}^{\mu}/p_{\mathbf{h}}\left(\mu\right)$ approximates the classification of our model with 98% fidelity. This behavior is reminiscent of K & H’s DAM [16], where only a few memories participate in the classification of each data point. The 1-nearest neighbor classifier approximates DAM classification more faithfully for correctly classified data points (approximately 99% fidelity) than incorrectly classified data points (approximately 70% fidelity).
@@
As before, we do so using constrained SGD (see Appendix C) with the initialization and learning rate described in Appendix D. Equivalently, we can also view this algorithm as minimizing the combined loss
@@
where $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x}^{*\mu}|\mathbf{w},\mathbf{p}\right)=\sum_{y=0}^{C}\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x}^{*\mu},y|\mathbf{w},\mathbf{p}\right)$ (see Eq. 14). $\mathcal{L}_{\text{cond}}\left(\mathbf{w},\mathbf{p}\right)$ is called the minimum entropy regularization term in unsupervised machine learning [73]. Intuitively, it encourages the DAM to learn different class weights $p^{\gamma}_{y}$ for each class $y$, which is not possible by minimizing only $\mathcal{L}_{\text{margin}}\left(\mathbf{w},\mathbf{p}\right)$ because $\mathcal{P}_{\beta,\varsigma}\left(\mathbf{x}^{*\mu}|\mathbf{w},\mathbf{p}\right)$ does not depend on $y$. Exploiting this characteristic, we train our DAM on patches of MNIST digits [50] and find that it learns reasonable latent classes $y=\operatorname*{argmax}_{y^{\prime}}\left\{p^{\mu}_{y^{\prime}}\right\}$ for the memories $\mathbf{w}^{\mu}$ (see Fig. 5 and Figs. 8, 9, 10 and 11 of Appendix J). This approach could potentially be useful for feature extraction [74].
@@ 4.3 Fast training with splitting steepest descent
Machine learning models that experience permutation symmetry breaking as described in Section 3.2—such as K & H’s DAM [24], RBMs with binary units [62, 63] and Gaussian mixtures [40, 41, 71] —have characteristic tree-shaped learning curves. Intuitively, permutation symmetry-breaking transitions are points where model parameters differentiate from each other, so they correspond to the bifurcations of the tree. In the left panel of Fig. (6), we show that the learning dynamics of our DAM also follows a tree of permutation symmetry-breaking transitions.
@@
The saddle-point hierarchy principle introduced in Section 3.2 suggests the following idea to accelerate learning: train a relatively narrow DAM for cheap, then repeatedly duplicate (or “split”) some of the hidden units, escape the corresponding saddle point and continue training. Intuitively, we expect to save a lot of computing resources by using relatively few hidden units at the start of training, when the learning dynamics follows a few branches close to the root of the learning dynamics tree. The splitting steepest descent algorithm introduced in [29] formalizes this idea in an efficient and theoretically motivated way. We implement a variant of this algorithm that takes the constraint $\mathbf{w}^{\mu}\in S^{N-1}$ into account, but is otherwise very similar to the original version (see Alg. 1), with fast splitting implemented as in [75]. The constraint $\mathbf{w}^{\mu}\in S^{N-1}$ only matters in steps 5 and 11, which are also arguably the most conceptually difficult parts of Alg. (1), so we explain them in detail in Appendix K. Fig. (6) shows that the learning dynamics tree of splitting steepest descent has a more sparsely populated trunk than that of SGD without splitting (see Fig. 6). In other words, by using a relatively small number $P$ of memories at the beginning of training, we reduce the total number of values $P\times T$ that they take over a period of time $T$, which is consistent with our intuition that it can save computational resources. The situation could have been different if using fewer memories slowed down training.
@@
Figure 6: Overlaps $m^{\mu_{*}\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)=\sum_{i=1}^{N}x^{*\mu_{*}}_{i}w^{\mu}_{i}$ between the first 1000 digits $\mathbf{x}^{*\mu_{*}}=\left\{x^{*\mu_{*}}_{i}\right\}_{i=1}^{N}$ of the MNIST training set [50] and the memories $\mathbf{w}^{\mu}=\left\{w^{\mu}_{i}\right\}_{i=1}^{N}$ of our dense associative memory (DAM) model while it is learning them. Each point is one of the high-dimensional magnetization vectors $m^{\cdot\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)=\left\{m^{\mu_{*}\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)\right\}_{\mu_{*}=1}^{P^{*}}$ projected onto a two-dimensional plane using the UMAP algorithm [76]. On the left, the $m^{\cdot\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)$ found during training without splitting steepest descent. On the right, the $m^{\cdot\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)$ found with splitting steepest descent (Alg. 1). In both cases, UMAP is trained on the $m^{\cdot\mu}\left(\mathbf{x}^{*},\mathbf{w}\right)$ found without splitting steepest descent. The classes $y=\operatorname*{argmax}_{y^{\prime}}\left\{p^{\mu}_{y^{\prime}}\right\}$ of the memories $\mathbf{w}^{\mu}$ are color coded in the legend. 7 and 1 are the most numerous classes, so they have the top two largest entries in $\sum_{\gamma=0}^{P}p^{\mu}_{y}=p_{\mathbf{q}}\left(y\right)$ (see Appendix D), which is why the memories are classified as either 7 or 1 at the beginning of training.
@@
14:return $\triangleright$ See Appendix K and [29, 75] for details about steps 5 and 11
We now establish a methodology that we will use to compare the training times of DAMs trained on MNIST [50] and Fashion-MNIST [77] using the effective loss (Eq. 12) with and without splitting steepest descent. It is more interesting and meaningful to compare the training times of NNs with similar performance. Therefore, we pick hyperparameters such that DAMs trained with and without splitting have similar classification accuracy. For that purpose, we find the general region of hyperparameter space where DAMs trained without splitting steepest descent have the best classification accuracy. The accuracy does not change much in this region, so we pick generic hyperparameters inside. Next, we manually tune the hyperparameters of splitting steepest descent so that the resulting DAMs have comparable accuracy. The accuracy obtained for Fashion-MNIST is generically slightly higher with splitting steepest descent than without it. In contrast, splitting steepest descent does not significantly affect the accuracy obtained for MNIST.